# WorldGUI: Dynamic Testing for Comprehensive Desktop GUI Automation

### v2

@@ Abstract
Current Graphical User Interface (GUI) agents have achieved outstanding performance in GUI element grounding. However, planning remains highly challenging, especially due to sensitivity to the initial state of the environment. Specifically, slight differences in the initial state—such as the target software not being open or the interface not being in its default state—often lead to planning errors. This issue is widespread in real application scenarios, but existing benchmarks fail to evaluate it. In this paper, we present WorldGUI, a novel GUI benchmark that designs GUI tasks with various initial states to simulate real computer-user interactions. The benchmark spans a wide range of tasks across 10 popular software applications, including PowerPoint, VSCode, and Adobe Acrobat. In addition, to address the challenges of dynamic GUI automation tasks, we propose GUI-Thinker, a holistic framework, leveraging a critique mechanism, that effectively manages the unpredictability and complexity of GUI interactions. Experimental results demonstrate that GUI-Thinker significantly outperforms Claude-3.5 (Computer Use) by 14.9% in success rate on WorldGUI tasks. This improvement underscores the effectiveness of our critical-thinking–based framework in enhancing GUI automation. The code is available at https://github.com/showlab/WorldGUI
Machine Learning, ICML
@@ 1 Introduction
Graphical User Interface (GUI) automation has become a prominent research area, driven by the need to enhance user productivity. This domain encompasses software usage, file management, office design, coding, and web browsing. Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o (OpenAI, 2023) and Claude-3.5 (Anthropic, 2024), GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to help the user.
Figure 1: The comparison of static and dynamic testing processes. Our WorldGUI takes the first step to facilitate comprehensive GUI evaluation with various initial states. The red node represents an incorrect state.
@@
GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition (He et al., 2016) and visual question answering (Antol et al., 2015). However, current GUI benchmarks such as OSWorld (Xie et al., 2024) and WebArena (Zhou et al., ) do not capture this dynamism. As shown on the left side of Fig. 1, most GUI benchmarks focus on initial and final states, measuring success rates but overlooking the changing initial conditions present in real GUI scenarios. These benchmarks often ignore situations where: (1) The software interface is not in its default state. (2) The agent might get user queries at any time. (3) Differences in agent robustness, where agents with the same low success rate (e.g. 20%) may vary in their ability to self-verify or self-correct, but these abilities are not measured in a static setting. As a result, these benchmarks fail to fully assess the capabilities of GUI agents.
In this paper, we take the first step toward comprehensive GUI agent testing by designing GUI tasks with various initial states. As illustrated on the right side of Fig. 1, the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default conditions, allowing tasks to start from intermediate states where users may seek assistance at any point (see Fig. 1 (b)). (2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing steps (see Fig. 1 (c)) or introducing new steps (see Fig. 1 (d)) to ensure task progression. By incorporating these elements, WorldGUI better mirrors real-world GUI environments, enabling a more accurate and thorough assessment of GUI agent capabilities.
Specifically, we present WorldGUI, a new benchmark featuring 10 desktop software applications and 315 tasks, including PowerPoint, Word, Excel, VSCode, etc. For each task, we provide a user query, an instructional video, and the corresponding project file. To ensure the task quality, we engaged four trained annotators skilled in using these applications and proficient in constructing data with scripts and agents for annotation. To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions. (See details in Sec. 3.)
In addition, we introduce a novel GUI agent framework, GUI-Thinker, which builds upon critical thinking philosophy, an aspect less emphasized in previous GUI agents (Hong et al., 2024; Cheng et al., 2024; Lai et al., 2024; Agashe et al., 2024; Wu et al., 2024). In dynamic GUI environments, application settings may not be in default configurations. This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy. Through our analysis of real-world GUI scenarios, we identify three critical designs for a comprehensive agent: (1) Post-Planning Critique, (2) Pre-Execution Validation, and (3) Post-Action Evaluation. Specifically, GUI-Thinker comprises five core components: Planner, Planner-Critic, Step-Check, Actor, and Actor-Critic. We argue that these components are fundamental for effective GUI agents.
To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in GUI automation and propose a new benchmark WorldGUI which designs the GUI tasks with various initial states to simulate the real interactions; (2) We introduce GUI-Thinker, a comprehensive GUI framework. GUI-Thinker incorporate the thinking into the overall agent design, which provides valuable insights and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling complex tasks.
2 Related Work
@@
Figure 2: WorldGUI: An illustration of our proposed real-world GUI benchmark. The left shows that for each task, WorldGUI provides a user query, instructional video, and pre-actions. The pre-actions lead to different initial states. The key characteristic of our WorldGUI is the various initial states of the same task to stimulate the real-world testing process. The right shows the software included in our benchmark and the interactions about testing the agents in our GUI environment.
2.1 GUI Benchmarks
GUI benchmarks are essential for evaluating the performance and robustness of GUI agents. For web applications, WebShop (Yao et al., 2022), and WebArena (Zhou et al., ) are two text-based GUI benchmarks, the GUI information is formatted in the text style which is limited to reflect the dynamic GUI state changes. In OS environments, OSWorld (Xie et al., 2024) is a comprehensive benchmark including various operating systems with real applications. Mobile benchmarks MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) propose two GUI benchmarks of mobile applications. Windows-related benchmarks like AssistGUI (Gao et al., 2024) and WindowAgentArena (Bonatti et al., 2024) propose a list of real tasks in the Windows platform. However, these benchmarks primarily rely on a static testing process and do not adequately capture the complexity and dynamic nature of GUI environments. As a result, they are insufficient for comprehensively evaluating GUI agents. We summarize the differences between WorldGUI and other related benchmarks in Tab. 5.
2.2 GUI Agents
Building upon MLLMs, GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to help the user. CogAgent (Hong et al., 2024) is a vision language model focused on GUI understanding to facilitate GUI navigation, while SeeClick (Cheng et al., 2024) and SeeAct (Zheng et al., 2024) focus on the GUI grounding for enhancing the task performance. MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) are proposed to design the agent on mobile device. Ferret-UI (You et al., 2025) is another representative work focusing on enhancing the grounding ability in the IOS platform. AutoGLM (Liu et al., 2024) propose the GUI modes capable of learning through autonomous environmental interactions by reinforcing existing GUI models. These agents have shown their ability in GUI understanding (e.g., GUI elements grounding) or action prediction but still face limitations in handling dynamic and complicated full GUI tasks. Therefore, to enhance GUI automation in dynamic environments, we propose GUI-Thinker that improves adaptability in complex GUI settings and enables agents to effectively handle unpredictable interface changes.
2.3 Critical Thinking in Agents
Recent advancements in foundation models and agents, particularly in LLMs such as OpenAI-o1 (OpenAI, 2024) and Deepseek-R1 (DeepSeek-AI et al., 2025), have increasingly incorporated thinking processes before providing answers to effectively handle challenging reasoning tasks. The LLM-based agents utilize verify-then-correct process to evaluate and refine intermediate reasoning steps or outputs, ensuring logical coherence and consistency. One notable LLM-based agent framework, Reflexion (Shinn et al., 2024), demonstrates the effectiveness of self-reflection in solving complex tasks. Furthermore, Critic (Gou et al., 2023) integrates external tools into the critique process, leveraging them to improve performance. Noticing the GUI task is lengthy and complicated, the verify-then-correct process is highly suitable for the GUI scenario. Which is not only aims to enhance the reasoning performance but also indispensable to designing the key module Actor-Critic (Konda & Tsitsiklis, 1999) to ensure task completion. A closely related work, AssistGUI (Gao et al., 2024), integrates a critical module only after the Actor module to evaluate action completion. Building upon it, we introduce two additional critical modules: Planner-Critic, applied after the Planner, and Step-Check, applied before the Actor. These two modules lead to a comprehensive and fundamental GUI agent framework GUI-Thinker which will provide insights for future GUI agent design.
@@ 3.1 Task Formulation
GUI Automation Definition. The GUI automation task can be considered a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R})$ with state space $\mathcal{S}$, observation $\mathcal{O}$, action space $\mathcal{A}$, transition function $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, and reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. In our setting, given a natural language query $q$, eg., Format the slide background with gradient fill that describes a specific task in high-level, along with an instructional video $v$ as a supplement that more detailed illustrates how to complete it, the agent first get the observation $o_{t}\in\mathcal{O}$ from the state $s_{t}\in\mathcal{S}$ in the execution environment and then generate the executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in\mathcal{S}$ and a new observation $o_{t+1}\in\mathcal{O}$. The process repeats until the task is finished or failed. The reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ here returns a binary integer at the final step indicating the task completion status.
@@
WorldGUI Task Definition. As illustrated in Fig. 2, each GUI task is paired with a user query and an instructional video. To achieve state diversity within each task, we generate various initial states that converge to the same final state, resulting in distinct ground truth (GT) plans for each case. This is accomplished through the use of pre-actions, which consist of a sequence of executable code to initialize tasks from different initial states. With the augmentation of initial states, WorldGUI is capable of mimicking the different testing scenarios as shown in Fig. 1.

### v3

# WorldGUI: An Interactive Benchmark for Desktop GUI Automation from Any Starting Point

@@ Abstract
GUI agents have achieved outstanding performance in GUI element grounding. However, planning remains highly challenging, especially due to the sensitivity to the initial state of the environment. Specifically, slight differences in the initial state-such as the target software not being open or the interface not being in its default state, often lead to planning errors. This issue is widespread in real application scenarios, but existing benchmarks fail to evaluate it. To address this gap, we introduce WorldGUI, a comprehensive GUI benchmark containing tasks across ten widely used desktop and web applications (e.g., PowerPoint, VSCode, Acrobat), each instantiated with diverse initial states to simulate authentic human–computer interactions. Complementing this, we propose WorldGUI-Agent, a universal framework that unifies three core modules: Planner-Critic for high-level plan refinement, Step-Check for intermediate verification, and Actor-Critic for action-level optimization to proactively detect and correct errors. Experimental evaluation shows that WorldGUI-Agent outperforms the outstanding existing model (Claude-3.5 Computer Use) by 12.4% in success rate on WorldGUI, and achieves a 31.2% overall success rate on WindowsAgentArena, surpassing the prior state-of-the-art by 11.7%. Our analysis further reveals that dynamic augmentation tasks and desktop environments pose substantial hurdles, underscoring the necessity of adaptive planning and feedback-driven execution for advancing real-world GUI automation. The code and data are available at https://github.com/showlab/WorldGUI.
00footnotetext: Corresponding author.
Figure 1: Software taxonomy of WorldGUI and the performance comparison of GUI agents. The left shows 5 main groups and 10 software in our WorldGUI. The right shows that WorldGUI-Agent surpasses previous SOTA GUI agents significantly.
@@ 1 Introduction
Graphical User Interface (GUI) automation has become a prominent research area, driven by the need to enhance user productivity. This domain encompasses software usage, file management, office design, coding, and web browsing. Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o gpt4o and Claude-3.5 claude3.5 , GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to enhance productivity efficiency.
@@
GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition he2016deep and visual question answering antol2015vqa ; VQAv2 . However, current online GUI benchmarks such as WebArena webarena , WebVoyager webvoyager , and WindowsAgentArena windowsagentarena do not capture this dynamism. Currently, most GUI benchmarks OSWorld ; windowsagentarena ; assistgui ; webarena ; visualwebarena ; webvoyager focus on initial and final states, measuring success rates but overlooking the state variety in real GUI scenarios. These benchmarks often ignore situations where: (1) The software interface is not in its default state. (2) The human-computer interactions may start from the intermediate state of a specific task. (3) Differences in agent robustness, where agents with the same low success rate (e.g., 20%) may vary in their ability to self-reflection, but these abilities cannot be measured in a static setting. As a result, these benchmarks fail to comprehensively assess the GUI agents.
@@
Figure 2: WorldGUI. The left shows that for each task, WorldGUI provides a user query, instructional video, and pre-actions. The pre-actions lead to different initial states. The key characteristic of our WorldGUI is the various initial states of the same task to stimulate the real-world testing process. The right shows the software included in our benchmark and the interactions about testing the agents in our GUI environment.
In this paper, we take the first step toward comprehensive GUI evaluation by designing GUI tasks with various initial states. We consider that the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default initial conditions, allowing tasks to start from intermediate states where users may seek assistance at any point. (2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing plans or introducing new steps to ensure task execution. By incorporating these situations into the benchmark design, WorldGUI better mirrors real-world GUI interactions, enabling a more accurate and thorough assessment of GUI agent capabilities. Specifically, WorldGUI embraces 10 widely-used desktop applications with 611 tasks in total. For each task, we create a user query, an instructional video, and the corresponding project file. We engaged four trained annotators skilled in using these applications for annotation. To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions.
In addition, we introduce a new GUI agent framework, WorldGUI-Agent, which builds upon critical thinking design principle, an aspect less emphasized in previous GUI agents hong2024cogagent ; cheng2024seeclick ; autowebglm ; AgentS ; osatlas . In dynamic GUI environments, application settings may not be in default configurations. This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy. Through our analysis of real-world GUI scenarios, we identify three design principles for GUI agents: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation. We argue that these components are fundamental and universal for GUI agents.
To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in the online GUI testing and propose a new benchmark called WorldGUI; (2) We introduce WorldGUI-Agent, a fundamental and universal GUI framework that incorporates critical thinking into the overall agent design, providing valuable insight and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling GUI tasks (see Figure 1).
@@ 2 WorldGUI Benchmark
Table 1: Comparison with other interactive GUI benchmarks. WorldGUI is a unique benchmark that embraces diverse initial states and better reflects the authentic interactions in GUI scenarios. Env?: Indicates whether an environment is required to be deployed.
Benchmarks Softwares Tasks Platform Env? Inst. Video? GT Plan Diverse Contextual
Init. State? Variability?
WebArena webarena 6 812 Web Yes
VisualWebArena visualwebarena 3 910 Web Yes
WebVoyager webvoyager 15 643 Web Yes
AutoDroid AutoDroid 13 158 Android OS Yes
AndroidWorld androidworld 20 116 Android OS Yes
AgentStudio zheng2025agentstudio 9 205 Desktop + Web Yes
Mobile-Eval wang2024mobile 10 30 Android OS Yes
APPAgent zhang2023appagent 10 50 Android OS Yes
OSWorld OSWorld 10 369 Desktop Yes
AssistGUI assistgui 9 100 Windows No
WindowAgentArena windowsagentarena 11 154 Windows Yes
WorldGUI 10 611 Win. + Web No
@@ 2.1 Task Formulation
GUI Automation Definition. The GUI automation task can be considered a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R})$ with state space $\mathcal{S}$, observation $\mathcal{O}$, action space $\mathcal{A}$, transition function $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, and reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. In our setting, given a natural language query $q$, eg., Format the slide background with gradient fill that describes a specific task in high-level, along with an instructional video $v$ as a supplement that more detailed illustrates how to complete it, the agent first get the observation $o_{t}\in\mathcal{O}$ from the state $s_{t}\in\mathcal{S}$ in the execution environment and then generate the executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in\mathcal{S}$ and a new observation $o_{t+1}\in\mathcal{O}$. The process repeats until the task is finished or fails. The reward function $\mathcal{R}$: $\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ here returns a binary integer at the final step ,indicating the task completion status.
@@
WorldGUI Task Definition. As illustrated in Figure  2, to achieve state diversity within each task, we generate various initial states that converge to the same final state, resulting in distinct ground truth (GT) plans for each case. This is accomplished through the use of pre-actions, which consist of a sequence of executable code to initialize tasks from different initial states. With the augmentation of initial states, WorldGUI is capable of mimicking the different testing scenarios. We additionally summarize the differences between WorldGUI and other close interactive benchmarks in Table 1.